Integrated Nanosystems Research Facility

MA6 is down

The INRF cleanroom will be closed tomorrow in order for Facilities to work on the compressor. BION will remain open.

Facilities will be shutting off power in that room for approximately 30min at 10am. Once the power is back on feel free to use the room but make sure the bench is functioning properly before starting your process. Please notify staff if not.

One of the exhaust fans that support those rooms is down and as a precautionary measure we are not allowing any work in those rooms. Facilities mentioned it should be back up by 12pm tomorrow.  We will let you know if its up sooner.

The E-Beam 1 is now back up. The Trion and Mark 50 are still down.

INRF Staff

From Facilities:

Update on the progress made since Wednesday and the plan moving forward

–          On Thursday NuAge and Plumbing sub cut open the wall on the third floor, installed a scaffolding system and replaced the defective parts. AGI was on site to address leaking issues and to relocate air movers and dehumidifiers. Tony Garrit (Central Plant) charged the line and the fitting was noted to be working properly now

–          On Friday with the leak being fixed, AGI set up containment at the end of affected corridors and around lab equipment to protect from dust as the space starts to dry out. AGI will do their best to cover and create a barrier around the equipment identified in Fridays walk through with Jordan Terrones and Jake Hes, however the facility understands that there is the possibility that dust will still get on equipment unless it is completely disconnected and removed from the space.

–          So far to note, floor tiles are going to need to be removed from W2310, W2311, W2314, W2315, W2318, W2320, & W2321 to see the extent of the water damage under the floor. As needed, flooring will be removed up to and around larger equipment and around table legs where items cannot be moved. Ceiling tiles will need to be removed in rooms W2310, W2312, & W2314 to see how far down the corridors the water travelled.

–          On Saturday AGI will have a crew of 5 guys total to remove damaged floor and ceiling tiles. Due to the extensiveness of the work, AGI may need to have a smaller crew to work on Sunday.

–          On Monday AGI will provide an assessment report of their findings and the extent of the damage.

–          After that time, the university can begin to plan the remediation of the space.

–          Central Plan was notified to shut down Eng. Gateway AHU’s 6 &7 that control cleanroom space.

–          AGI brought in Negative Air Machines to limit the dust as the space begins to dry out.

–          On Saturday AGI finished installing protection around lab equipment and began removing ceiling tiles and floor tiles from affected areas. As we began pulling up tiles the extent of how far the water travelled was realized and it was determined that AGI would need to continue work on Sunday.

–          On Sunday AGI continues to remove flooring tiles

–          Monday will consist of AGI wiping down the space of surface dust and providing assessment of their findings.

As of now the cleanroom is still closed and we will keep you updated if things change. The BION has not been effected and will remain open.

Yesterday we experienced a flood in the first half of the cleanroom due to a break in the chill-water system. Thankfully it happened where none of the critical equipment were located. Facilities is working to clean up the water and bring power back to the cleanroom. Due to the extent of the damage we are expecting the lab to be down for the rest of the week and our hope is to be up and running early next week. We will keep you posted.

In other news, the compressed air system should be operational now (once the power is restored) and parts for the Trion, Mark 50 and Ebeam1 have arrived.

NOTE: BION was unaffected and will remain open.

Due to a flood in the cleanroom the INRF will be closed for the remainder of the day for cleanup. The BION will remain open.

Curtailment of power to electrical outlets for construction tie-in of new power circuits. Jake will start shutting down equipment Thursday Night (tomorrow 04/07) and will start bringing them back online Friday after they bring the power back. We will make updates as the tools become available.

Friday April 8, 2016, 5:00am – 8:00am

BION Power Outage

This weekend we experienced that the compressed air system has not been able to run two consecutive days without an alarm.  The alarm we have been experiencing is low oil pressure / discharge temperature and high inter stage temperature. We are working with Facilities and the contractor to address these issues.  This means that quite a few pieces of equipment are still shut down due to the CDA. We share your frustration and apologize for the inconvenience. We will keep you updated.